# The Fall of an Algorithm: Characterizing the Dynamics Toward Abandonment

### v1

@@ Appendix B Methodology (extended)
To identify additional cases that met our criteria outside of the U.S., we conducted additional keyword searches on Google of names of regions/countries and our keywords from before, e.g., “China algorithmic harm”, etc.
B.1. Thematic analysis (continued)
@@
For illustrative purposes, we walk-through how we coded an event timeline for an example case: the National Eating Disorder Association (NEDA)’s Tessa Wellness Bot (see the final timeline in Section C below). We first learned of this case by searching the AI Incident Database (McGregor, 2020). We began learning about the case by reading the news articles linked on its AIID webpage. While reading these articles, we closed-coded case metadata, and also noted any other important events that appeared to be relevant to the algorithmic harm(s) and decisions made by the model owners. For example, we learned from a Vice article (Xiang, 2023) that shortly before the Tessa bot came under criticism for its potential to inflict harm, a news story broke that NEDA had decided to fire all of its hotline staffers following their decision to unionize, and intended to “replace them” by instead directing people to Tessa, a chatbot that had been in operation since February 2020. From this article, we also learned the name of the organization that developed Tessa, and gained insight into Tessa’s design (e.g., how users interacted with Tessa online). Two more recent articles linked on the AIID webpage (Wells, 2023; Hoover, 2023) interviewed users who took to social media to criticize the prominent NEDA firing and the Tessa chatbot. These articles directly linked to social media posts where we could directly engage with users’ critiques and discussions about the algorithm, such as Sharon Maxwell’s Instagram posts (link). We referenced these interviews and critiques to better understand the nature of the algorithm’s potential for harm and who might be adversely impacted. Finally, we also referred to these articles and primary sources (e.g., NEDA’s Instagram announcement, link) to understand the organization’s decision to abandon the algorithm. We conducted a final confirmatory Google search to check if the Tessa Bot (or any similar service created by Cass) is still in operation today.
Appendix C Event Timeline Figure (Expanded)
@@
February 2017, release. Pittsburgh begins a hotspot-based predictive policing pilot with researchers at CMU (through ”Metro21”). They do not notify or involve the public in the development of the program.
@@
May 2023, discovery of potential harm. NPR runs an exposé on how NEDA fired all helpline staff, and intends to ”replace them” with Tessa.
May 2023. Five days after the NPR exposé, fat-positive activist Sharon Maxwell shares screenshots of harmful interactions with Tessa, in which the bot gave advice that might trigger an eating disorder (e.g., to count calories and lose weight). Her post goes viral (dissemination). A NEDA VP comments on Sharon’s post that she is lying and asks Sharon to provide her with more evidence that the interaction actually occurred (dialogue).

### v2

@@ Appendix A Methodology (extended)
To identify additional cases that met our criteria outside of the U.S., we conducted additional keyword searches on Google of names of regions/countries and our keywords from before, such as “China algorithmic harm”.
A.1. Thematic analysis (continued)
@@
For illustrative purposes, we walk-through how we coded an event timeline for an example case: the National Eating Disorder Association (NEDA)’s Tessa Wellness Bot (see the final timeline in Section B below). We became aware of this case by searching the AI Incident Database (McGregor, 2020). We began learning about the case by reading the news articles linked on its AIID webpage. While reading these articles, we closed-coded case metadata, and also noted any other important events that appeared to be relevant to the algorithmic harm(s) and decisions made by the model owners. For example, we learned from a Vice article (Xiang, 2023) that shortly before the Tessa bot came under criticism for its potential to inflict harm, a news story broke that NEDA had decided to fire all of its hotline staffers following their decision to unionize, and intended to “replace them” by instead directing people to Tessa, a chatbot that had been in operation since February 2022. From this article, we also learned the name of the organization that developed Tessa, and gained insight into Tessa’s design (e.g., how users interacted with Tessa online). Two more recent articles linked on the AIID webpage (Wells, 2023; Hoover, 2023) interviewed users who took to social media to criticize the prominent NEDA firing and the Tessa chatbot. These articles directly linked to social media posts where we could directly engage with users’ critiques and discussions about the algorithm, such as Instagram posts (link). We referenced these interviews and critiques to better understand the nature of the algorithm’s potential for harm and who might be adversely impacted. Finally, we also referred to these articles and primary sources (e.g., NEDA’s Instagram announcement, link) to understand the organization’s decision to abandon the algorithm. We conducted a final confirmatory Google search to check if the Tessa Bot (or any similar service created by Cass) is still in operation today.
Appendix B Event Timeline Figure (Expanded)
@@
February 2017, release. Pittsburgh begins a hotspot-based predictive policing pilot with researchers at CMU (through the Metro21 Smart Cities Institute). They do not notify or involve the public in the development of the program.
@@
May 2023, discovery of potential harm. NPR runs an exposé on how NEDA fired all helpline staff following their decision to unionize, and intends to ”replace them” with Tessa.
May 2023. Five days after the NPR exposé, a fat-positive activist shares screenshots of harmful interactions with Tessa, in which the bot gave advice that might trigger an eating disorder (e.g., to count calories and lose weight). Her post goes viral (dissemination). A NEDA VP comments on the author’s post that she is lying and asks the author to provide her with more evidence that the interaction actually occurred (dialogue).
@@
Appendix C Case List
Interact with and download our case dataset online at this (link)!
| Case Name | Case Description | |
| Online platforms | ||
| 1 | Twitter Cropping Algorithm (Yee et al., 2021) | Users hypothesized that an algorithm that cropped images to show feed previews compatible with device aspect ratios was more likely to crop out Black people (or other minority groups, like women) in images with multiple people. |
| Chatbot | ||
| 4 | U.S. National Eating Disorder Assocation (NEDA)’s “Tessa” Wellness Bot (Wells, 2023) | Shortly NEDA fired all of their hotline staff (following their decision to unionize), activists posted on social media about how Tessa (NEDA’s chatbot) gave harmful and triggering advice. |
| 5 | Meta’s Galactica (gal, 2022) | Within three days of releasing a public demo of their Galactica language model, Meta took down the public demo and re-branded Galactica as ”available for researchers” only. |
| 6 | Scatter Lab’s ”Lee-Luda” Chatbot (https://www.theguardian.com/profile/justinmccurry, [n. d.]) | Seoul-based startup’s AI-chatbot ”Lee-Luda” was removed from Facebook messenger 20 days upon launch after users complained it was using offensive language, discriminatory and generated hate speech towards LGBTQ folks and people with disabilities. |
| Employment | ||
| 7 | Amazon AI Recuriting Tool (Dastin, [n. d.]) | Amazon disbanded an internal team whose goal was to train ML to screen resumes due to algorithmic gender bias. |
| Child Welfare | ||
| 8 | Oregon DHS’s Safety at Screening Tool (ap2, 2022) | Oregon DHS decided to stop using an AI-based screening tool (which is modeled after the AFST), in the wake of an April AP news article reporting that these tools have the ability to ”heighten racial disparities in the child welfare system”. |
| Case Name | Case Description | |
| Child Welfare | ||
| 11 | New Zealand’s ”Vulnerable Children PRM” (Predictive Risk Modeling) (Bukowitz and O’Loughlin, 2022) | New Zealand’s Social Development Minister intervened and halted the deployment of the algorithm after seeing major concerns from additional research during the request for its ethics approval for further study. |
| Family & Social Services | ||
| 12 |
Teen Pregnancy Risk
Prediction in Argentina (Aplicada, 2018) |
| 13 |
The Spanish Ministry of the Interior’s “VioGén”
Gender Violence Recidivism Risk
Assessment (eti, 2022) |
The Eticas Foundation partnered with victims to publish an external audit after the Spanish Ministry declined requests of auditing the algorithm that underestimated level of risk in domestic and intimate partner violence cases, leading to adverse outcomes. |
| Benefits | ||
| Case Name | Case Description | |
| Benefits | ||
| 18 |
Netherlands’ System Risk
Indication (SyRI) (van Veen, 2020) |
Local civil rights organizations filed a case against the state to stop using an algorithm to detect welfare benefits fraud. A Dutch court ruled that the algorithm violated the European Convention on Human Rights (ECHR) protections and halted its operation. |
| Healthcare | ||
| 19 | EPIC’ Sepsis Prediction Tool (Ross, 2021) | Epic, an American healthcare software company (and prominent EHR vendor), launched the Epic Sepsis Model (ESM) in 2017, a proprietary linear model that alerts patients at high risk of sepsis. From 2021 onward, external and independent audits or investigations showed that the model was flawed; e.g., it had lower AUC than reported, had a high false alarm rate, etc. |
| 20 | US HHS and UNOS (the United Network for Organ Sharing)’s ”Acuity Circles Policy” (was, 2023) | A joint investigation by the Washington Post and The Markup exposed how a new shift in the UNOS allocation algorithm’s criteria in removing distance as a factor for organ donation and prioritizing sickest patients regardless of location, drastically decreased the number of transplant surgeries and increased deaths in poorer states. |
| Case Name | Case Description | |
| Property Tech | ||
| 24 | RealPage’s YieldStar Algorithm (Vogell et al., 2022) | RealPage, a Texas-based company that provides property management software for landlords, developed an algorithm, YieldStar, that suggests daily prices for open rental units. Critics claim that the software brings anti-trust concerns and has resulted in rent inflation across the U.S. |
| Education | ||
| 26 |
SAS Institute’s Educational Value-Added Assessment
System (EVAAS) in Houston (Capo and Bass, 2017) |
| Facial Recognition | ||
| 31 |
Facial Recognition in NYC
Atlantic Plaza Towers (Gagne, 2019) |
| Case Name | Case Description | |
| Facial Recognition | ||
| 32 | RiteAid’s Enrollement Database for Surveillance Systems (Dastin, 2020) | RiteAid, a large U.S. drugstore chain, deployed facial recognition surveillance systems in over 200 stores. Employees used the FR to surveil customers (e.g., ”persons of interest” who might attempt to engage in criminal activity). RiteAid stopped using FR after a Reuters investigation broke in 2020, and in 2023 were issued a five-year ban on their use of FR by the FTC. |
| 37 |
Uganda’s
”Intelligent Transport Monitoring System” (Nyeko, 2023) |
| Predictive Policing | ||
| 39 | City of Pittsburgh (in Pittsburgh, 2020a) | Pittsburgh City Council voted to ban ”obtaining, retaining, accessing, or using FR or predictive policing technology” after facing criticism from student activists for a predictive policing experiment in collaboration with researchers at Carnegie Mellon (Metro21). |